### v2

@@ 7. Motives of generically quasi-split varieties
Let $b_{1},\ldots,b_{r}$ be a homogeneous $\mathbb{F}_{p}$-basis of $\overline{\mathop{\mathrm{CH}}\nolimits}_{K}(Y)$ and let ${b^{*}_{1},\ldots,b^{*}_{r}\in\mathop{\mathrm{CH}}\nolimits_{K}(\overline{Y})}$ be a family of cycles dual to this basis with respect to the bilinear form ${\mathop{\mathrm{CH}}\nolimits_{K}(\overline{Y})\times\mathop{\mathrm{CH}}\nolimits_{K}(\overline{Y})\rightarrow\mathbb{F}_{p}}$, $(x,y)\mapsto\deg(x\cdot y)$.
@@
Let $i\in[1,r]$ and let $x_{i}\in\mathop{\mathrm{CH}}\nolimits_{K}(Y)$ be a preimage of $b_{i}\in\mathop{\mathrm{CH}}\nolimits_{K}(\overline{Y})$. We construct two correspondences $\alpha_{i},\beta_{i}\in\mathop{\mathrm{CH}}\nolimits_{K}(Y\times Y)$ as follows. We set $\beta_{i}=(x_{i}\times 1)\cdot\pi$ and define $\alpha_{i}$ as a cycle whose image under the surjection $\mathop{\mathrm{CH}}\nolimits_{K}(Y\times Y)\,\,-\!\!\!\twoheadrightarrow\mathop{\mathrm{CH}}\nolimits_{K}(\operatorname{Spec}F(Y)\times Y)\simeq\mathop{\mathrm{CH}}\nolimits_{K}(\overline{Y})$ is $b_{i}^{*}$. Note that by construction of $\alpha_{i}$ and $\beta_{i}$ we have $\operatorname{mult}(\beta_{i}\circ\alpha_{i})=1$ and, hence, $\operatorname{mult}(\pi\circ\beta_{i}\circ\alpha_{i}\circ\pi)=1$. By Corollary 7.3 we obtain $\overline{\pi\circ\beta_{i}\circ\alpha_{i}\circ\pi}=\overline{\pi}$. It follows that the correspondence $\overline{\alpha_{i}\circ\pi\circ\beta_{i}}\in\mathop{\mathrm{CH}}\nolimits_{K}(\overline{Y}\times\overline{Y})$ is a projector, which we denote by $\overline{\pi}_{i}$. By construction of the correspondences $\alpha_{i}$ and $\beta_{j}$, we have $\operatorname{mult}(\beta_{j}\circ\alpha_{i})=0$ if $i\neq j$. Thus, it follows from Corollary 7.3 that the projectors $\overline{\pi}_{1},\dots,\overline{\pi}_{r}$ are pairwise orthogonal.
@@
The above decomposition of the diagonal class $\Delta_{\overline{Y}}$ implies the decomposition ${\mathop{\mathrm{CH}}\nolimits_{K}(\overline{Y})=\operatornamewithlimits{\textstyle\bigoplus}_{i=1}^{r}\mathop{\mathrm{CH}}\nolimits_{K}(\overline{Y},\overline{\pi}_{i})}$. We claim that $\mathop{\mathrm{CH}}\nolimits_{K}(\overline{Y},\overline{\pi}_{i})=b_{i}\mathop{\mathrm{CH}}\nolimits_{K}(\overline{Y,\pi})$. Since ${\overline{\pi}^{*}_{i}=\overline{\beta}_{i}^{*}\circ\overline{\pi}^{*}\circ\overline{\alpha}_{i}^{*}}$ we have
@@
where the second inclusion follows from the equality $\overline{\beta}_{i}=(b_{i}\times 1)\overline{\pi}$. Moreover, we have an isomorphism of split motives $(\overline{Y},\overline{\pi}_{i})\simeq\overline{(Y,\pi)}(k)$ with $k=\operatorname{codim}b_{i}$. It follows that $\dim_{\mathbb{F}_{p}}\mathop{\mathrm{CH}}\nolimits_{K}(\overline{Y},\overline{\pi}_{i})=\dim_{\mathbb{F}_{p}}\mathop{\mathrm{CH}}\nolimits_{K}(\overline{Y,\pi})$. Therefore, the inclusion ${\mathop{\mathrm{CH}}\nolimits_{K}(\overline{Y},\overline{\pi}_{i})\subset b_{i}\mathop{\mathrm{CH}}\nolimits_{K}(\overline{Y,\pi})}$ is an equality and we have $\mathop{\mathrm{CH}}\nolimits_{K}(\overline{Y})=\operatornamewithlimits{\textstyle\bigoplus}_{i=1}^{r}b_{i}\mathop{\mathrm{CH}}\nolimits_{K}(\overline{Y,\pi})$. ∎
@@
Note that the projectors $\pi_{i}$ in the proof of the above lemma are $F$-rational. Therefore, the decomposition of the diagonal class $\Delta_{\overline{Y}}=\overline{\pi}_{1}+\dots+\overline{\pi}_{r}$ provides the motivic decomposition $M_{K}(Y)=\operatornamewithlimits{\textstyle\bigoplus}_{i=1}^{r}\mathcal{U}_{K}(Y)(\operatorname{codim}b_{i})$.
@@
[…]
@@
[…]
@@ 8. $J$-invariant for groups of outer type
[…]
@@
The respective quasi-split group $G^{qs}$ from sequence (5.3) is a central product of a group of type ${{}^{2}\mathrm{A}_{2n+1}}$ and a group of type (${}^{2}(2\,{{}^{1}\mathrm{A}_{2n+1}})$. Since the normed Chow motive of a generic maximal unitary Grassmannian is indecomposable, by Proposition 7.7 its normed Chow group over a splitting field does not contain rational cycles of positive codimension. Note also that $\mathop{\mathrm{CH}}\nolimits^{*}_{K}(R_{K/F}(\mathbb{P}^{2n+1}))=\mathbb{F}_{2}[R(h)]/(R(h)^{2n+2})$. The proof of Proposition 8.2 below shows that the subring of rational cycles on $X\times R_{K/F}(\mathop{\mathrm{SB}}(B))$ in the generic case is generated by the cycle $1\times R(h)^{j}$ with $j=2^{k}$ such that $2^{k}\parallel\frac{2n+2}{2}$ and $\operatorname{codim}R(h)=2$.
@@
Let $\mathfrak{X}$ be a generically quasi-split variety under the action of the group $\operatorname{Aut}_{K}(B,\tau)$. Let $U$ be the normed upper motive of $\mathfrak{X}$. Consider the variety ${\mathcal{Y}:=\mathfrak{X}\times R_{K/F}(\mathop{\mathrm{SB}}(A))}$. The variety $\mathcal{Y}$ is generically quasi-split under the action of the group ${\operatorname{Aut}_{K}(B,\tau)\times R_{K/F}(\operatorname{\mathrm{PGL}}_{1}(A))}$. Since both varieties $\mathfrak{X}_{F(\mathcal{Y})}$ and $\mathcal{Y}_{F(\mathfrak{X})}$ have rational points, the normed upper motives of $\mathfrak{X}$ and $\mathcal{Y}$ are isomorphic: $\mathcal{U}_{K}(\mathcal{Y})\simeq\mathcal{U}_{K}(\mathfrak{X})=:U$. Denote by $m$ and $F_{A}$ respectively the degree of $A$ and the function field of the variety $R_{K/F}(\mathop{\mathrm{SB}}(A))$.
@@
Our next goal is to describe the quotient $P(\overline{\mathop{\mathrm{CH}}\nolimits}_{K}^{*}(\mathcal{Y}),t)/P(\overline{\mathop{\mathrm{CH}}\nolimits}_{K}^{*}(\mathfrak{X}_{F_{A}}),t)$. We follow similar arguments as in [Zh22, §3]. We refer to Section 3 for the notation $R(h)$ below.
@@
Since both varieties $\mathfrak{X}$ and $R_{K/F}(\mathop{\mathrm{SB}}(A))$ are quasi-split over the function field of $\mathfrak{X}$, we can follow the proof of [Zh22, Lemma 3.2] by replacing the usual Chow rings $\mathop{\mathrm{CH}}\nolimits$ by the normed Chow rings $\mathop{\mathrm{CH}}\nolimits_{K}$. ∎
Using Lemma 8.6, similarly as in [Zh22, Proposition 3.4] we can show that the polynomial $P(\overline{\mathop{\mathrm{CH}}\nolimits}_{K}^{*}(\mathfrak{X}_{F_{A}}),t)$ divides $P(\overline{\mathop{\mathrm{CH}}\nolimits}_{K}^{*}(\mathcal{Y}),t)$ and the quotient is given by $F(t):=\operatornamewithlimits{\textstyle\sum}_{i\in J}t^{i}$, where
(8.7) $J=\{0\leq j\leq m\mid 1\times R(h)^{j}\text{ is }F\text{-rational in }\mathop{\mathrm{CH}}\nolimits_{K}^{*}(\overline{\mathcal{Y}})\}\,.$
@@
We show next that the parameter $k_{1}$ of the $J$-invariant satisfies $k_{1}<v_{2}(\deg B)$. We assume that we are in type ${}^{2}\mathrm{A}_{2n+1}$ and, thus, $\deg B=2n+2$. Consider the generically quasi-split variety $Y$ of parabolic subgroups of the group $\operatorname{Aut}_{K}(B,\tau)$ of type $\{1,2,\ldots,2n+1\}\setminus\{1,n+1,2n+1\}$. This variety is generically quasi-split by [Ti66, p. 55, type ${}^{2}\mathrm{A}_{n}$]. Then the normed Poincaré polynomial of $Y$ must be divisible by the Poincaré polynomial of the normed upper motive $U$ of the variety of Borel subgroups. Thus, the polynomial
@@
Denote $v_{2}(\mathop{\mathrm{ind}}A)$ by $i_{A}$. We claim that $j_{1}=\min\{k_{1},i_{A}\}$ and $k_{1}=v_{2}(\deg B)-1$. We know that $\min\{k_{1},i_{A}\}$ is an upper bound for $j_{1}$, hence, it is sufficient to show that $j_{1}\geq\min\{k_{1},i_{A}\}$. Recall that $j=2^{j_{1}}=\min\,J\backslash\{0\}$. It follows from Definition 8.7 of the set $J$ that the cycle $1\times R(h)^{j}\in\mathop{\mathrm{CH}}\nolimits_{K}^{*}(\overline{\mathfrak{X}\times R_{K/F}(\mathop{\mathrm{SB}}(A))})$ is $F$-rational. Then the cycle $1\times 1\times R(h)^{j}\in\mathop{\mathrm{CH}}\nolimits_{K}^{*}(\overline{\mathfrak{X}\times\mathcal{Z}\times R_{K/F}(\mathop{\mathrm{SB}}(A)))}$ is also $F$-rational, where $\mathcal{Z}$ is the variety of maximal parabolic subgroups of $\operatorname{Aut}_{K}(B,\tau)$ of type $\{1,2,\ldots,2n+1\}\setminus\{n+1\}$, i.e., the variety of totally isotropic ideals in $B$ of reduced dimension $n+1$.
Note that the variety $\mathfrak{X}$ possesses a rational point over the function field of the variety $\mathcal{Z}\times R_{K/F}(\mathop{\mathrm{SB}}(A))$. Therefore, by [Zh22, Lemma 2.3] the cycle
@@
is $F$-rational. It follows that ${R(h)^{j}\in\mathop{\mathrm{CH}}\nolimits_{K}^{*}(\overline{R_{K/F}(\mathop{\mathrm{SB}}(A))})}$ is rational over $F(\mathcal{Z})$. Therefore, passing to $K$ we find a rational cycle over $K(\mathcal{Z})$ of the form ${h^{j}\times h^{j}+\alpha\in\mathop{\mathrm{Ch}}\nolimits^{*}(\overline{\mathop{\mathrm{SB}}(A)\times\mathop{\mathrm{SB}}(A)})}$ in the ordinary Chow group, where $\alpha$ is some cycle which goes to zero in the respective normed Chow group.
@@
In the notation of Proposition 8.2 ($\deg B$ is even) we have: $j_{2}=0$ or $1$, and $j_{2}=0$ iff the discriminant algebra $D(B,\tau)$ (see [Ta76], [KMRT, Definition (10.28)]) is split.
@@
In the notation of Proposition 8.2 ($\deg B=2n+2$ is even) let $(j_{1},j_{2},j_{3},\ldots,j_{n+2})$ be the $J$-invariant of $(B,\tau)$. Then the $J$-invariant of $(B,\tau)_{F(\mathop{\mathrm{SB}}(D(B,\tau)))}$ equals $(j_{1},0,j_{3},\ldots,j_{n+2})$, i.e., apart from the value of $j_{2}$ the $J$-invariant does not change.
@@
The proof is the same as of [Zh22, Theorem 4.1(1)]. ∎
@@
By [Ka12a, Theorem 5.1] the ordinary Chow motive modulo $2$ of the generic orthogonal Grassmannian $\mathrm{OGr}$ of isotropic $(n-1)$-dimensional subspaces (it corresponds to the parabolic subgroup of type $\{1,2,\ldots,n-2\}$) is indecomposable and the same method of the proof shows that its normed motive is indecomposable as well. Since for a split algebra $A$ the upper motive of this Grassmannian is isomorphic to the upper motive of the variety of Borel subgroups, its Poincaré polynomial equals the Poincaré polynomial of ${}^{2}\mathrm{D}_{n}/P_{\{1,2,\ldots,n-2\}}$. The formula for the latter Poincaré polynomial follows directly from Proposition 6.1, namely, it equals
@@
Since the motive of the generic orthogonal Grassmannian of isotropic $(n-1)$-dimensional subspaces is indecomposable, it follows that the ring $C^{*}$ from sequence (5.3) equals in the ${}^{2}\mathrm{D}_{n}$ case the normed Chow ring $\mathop{\mathrm{CH}}\nolimits^{*}_{K}(\mathrm{OGr}(n-1,q))$ for a quasi-split quadratic form $q$ of dimension $2n$.
@@
The (normed) $J$-invariant of a central simple algebra with an orthogonal involution $(A,\sigma)$ of a non-trivial discriminant is the concatenation of the $J$-invariant of the quadratic form associated with $\sigma_{F(\mathop{\mathrm{SB}}(A))}$ (computed above) and ${j_{1}=\min\{k_{1},v_{2}(\mathop{\mathrm{ind}}A)\}}$.
@@
Let $U$ denote the normed upper motive of a generically quasi-split variety for the group $\operatorname{\mathrm{PGO}}^{+}(A,\sigma)$ and $U^{F_{A}}$ the normed upper motive of the same variety over the function field $F_{A}$ of the Severi–Brauer variety $\mathop{\mathrm{SB}}(A)$. Proceeding as in [Zh22, §3] we arrive at the following identity for the Poincaré polynomials
@@
We show now that $12^{k_{1}}\mid(\frac{\deg A}{2}-1)=n-$ 1. Consider the generically quasi-split variety $Y$ of parabolic subgroups of type $\{1,2,\ldots,n\}\setminus\{1,n-1,n\}$. This variety is generically quasi-split by [Ti66, p. 57, type ${}^{2}\mathrm{D}_{n}$]. Then the normed Poincaré polynomial of $Y$ must be divisible by the Poincaré polynomial of the normed upper motive $U$ of the variety of Borel subgroups. Thus, the polynomial
@@
Computing the quotient $\frac{P(\mathop{\mathrm{CH}}\nolimits^{*}_{K}(Y),t)}{P(\mathop{\mathrm{CH}}\nolimits^{*}_{K}(U),t)}$ we obtain $\frac{t^{n-1}-1}{t^{2^{k_{1}}}-1}$ which is a polynomial iff $2^{k_{1}}$ divides $n-1$.
@@
The normed $J$-invariant of the respective group of type ${}^{2}\mathrm{D}$ equals $(j_{1},j_{2},\ldots,j_{r})$, where $j_{i}$ of degree $d_{i}$ equals $1$, if $d_{i}=\dim\pi_{s-1}/2-1$ and equals 00 otherwise.
@@
where ``"``"`$\ldots$"` ` … " stands for a sum of monomials of degree $\geq 3$.
Since $A$ is split, note that the generator $w$ of $\mathop{\mathrm{CH}}\nolimits^{1}_{K}(\overline{X})\simeq\mathbb{F}_{2}$ is $F$-rational, and moreover $w^{2}\neq 0$. Hence, we have $P(\overline{\mathop{\mathrm{CH}}\nolimits}_{K}^{*}(X),t)=1+t+t^{2}+\ldots\,$, where ``"``"`$\ldots$"` ` … " stands for a sum of monomials of degree $\geq 3$.

### v3

@@ 7. Motives of generically quasi-split varieties
Let $b_{1},\ldots,b_{r}$ be a homogeneous $\mathbb{F}_{p}$-basis of $\overline{\mathop{\mathrm{CH}}\nolimits}_{K}(Y)$ and let $b^{*}_{1},\ldots,b^{*}_{r}\in\mathop{\mathrm{CH}}\nolimits_{K}(\overline{Y})$} be a family of cycles dual to this basis with respect to the bilinear form
$\mathop{\mathrm{CH}}\nolimits_{K}(\overline{Y})\times\mathop{\mathrm{CH}}\nolimits_{K}(\overline{Y})$ $\rightarrow\mathbb{F}_{p}$
$(x,y)$ $\mapsto\deg(x\cdot y)$
@@
Let $i\in[1,r]$ and let $x_{i}\in\mathop{\mathrm{CH}}\nolimits_{K}(Y)$ be a preimage of $b_{i}\in\mathop{\mathrm{CH}}\nolimits_{K}(\overline{Y})$. We construct two correspondences $\alpha_{i},\beta_{i}\in\mathop{\mathrm{CH}}\nolimits_{K}(Y\times Y)$ as follows. We set $\beta_{i}=(x_{i}\times 1)\cdot\pi$ and define $\alpha_{i}$ as a cycle whose image under the surjection $\mathop{\mathrm{CH}}\nolimits_{K}(Y\times Y)\,\,-\!\!\!\twoheadrightarrow\mathop{\mathrm{CH}}\nolimits_{K}(\operatorname{Spec}F(Y)\times Y)\simeq\mathop{\mathrm{CH}}\nolimits_{K}(\overline{Y})$ is $b_{i}^{*}$. Note that by construction of $\alpha_{i}$ and $\beta_{i}$ we have $\operatorname{mult}(\beta_{i}\circ\alpha_{i})=1$ and, hence, $\operatorname{mult}(\pi\circ\beta_{i}\circ\alpha_{i}\circ\pi)=1$. By Corollary 7.3 we obtain $\overline{\pi\circ\beta_{i}\circ\alpha_{i}\circ\pi}=\overline{\pi}$. It follows that the correspondence $\overline{\alpha_{i}\circ\pi\circ\beta_{i}}\in\mathop{\mathrm{CH}}\nolimits_{K}(\overline{Y}\times\overline{Y})$ is a projector, which we denote by $\overline{\pi}_{i}$. Moreover, $\overline{\pi}_{i}$ are rational, since so are the cycles $\overline{\alpha}_{i}$, $\overline{\pi}$ and $\overline{\beta}_{i}$. By construction of the correspondences $\alpha_{i}$ and $\beta_{j}$, we have $\operatorname{mult}(\beta_{j}\circ\alpha_{i})=0$ if $i\neq j$. Thus, it follows from Corollary 7.3 that the projectors $\overline{\pi}_{1},\dots,\overline{\pi}_{r}$ are pairwise orthogonal.
@@
The above decomposition of the diagonal class $\Delta_{\overline{Y}}$ implies the decomposition
$\mathop{\mathrm{CH}}\nolimits_{K}(\overline{Y})=\operatornamewithlimits{\textstyle\bigoplus}_{i=1}^{r}\mathop{\mathrm{CH}}\nolimits_{K}(\overline{Y},\overline{\pi}_{i})$}.
We claim that $\mathop{\mathrm{CH}}\nolimits_{K}(\overline{Y},\overline{\pi}_{i})=b_{i}\mathop{\mathrm{CH}}\nolimits_{K}(\overline{Y,\pi})$. Since $\overline{\pi}^{*}_{i}=\overline{\beta}_{i}^{*}\circ\overline{\pi}^{*}\circ\overline{\alpha}_{i}^{*}$}, where the upper star denotes the co-realization (see Section 3), we have
@@
where the second inclusion follows from the equality $\overline{\beta}_{i}=(b_{i}\times 1)\overline{\pi}$. Moreover, we have an isomorphism of split motives $(\overline{Y},\overline{\pi}_{i})\simeq\overline{(Y,\pi)}(k)$ with $k=\operatorname{codim}b_{i}$. It follows that $\dim_{\mathbb{F}_{p}}\mathop{\mathrm{CH}}\nolimits_{K}(\overline{Y},\overline{\pi}_{i})=\dim_{\mathbb{F}_{p}}\mathop{\mathrm{CH}}\nolimits_{K}(\overline{Y,\pi})$. Therefore, the inclusion
$\mathop{\mathrm{CH}}\nolimits_{K}(\overline{Y},\overline{\pi}_{i})\subset b_{i}\mathop{\mathrm{CH}}\nolimits_{K}(\overline{Y,\pi})$}
is an equality and we have $\mathop{\mathrm{CH}}\nolimits_{K}(\overline{Y})=\operatornamewithlimits{\textstyle\bigoplus}_{i=1}^{r}b_{i}\mathop{\mathrm{CH}}\nolimits_{K}(\overline{Y,\pi})$. ∎
@@
Note that the projectors $\overline{\pi}_{i}$ in the proof of the above lemma are $F$-rational. Therefore, the decomposition of the diagonal class $\Delta_{\overline{Y}}=\overline{\pi}_{1}+\dots+\overline{\pi}_{r}$ provides the motivic decomposition $M_{K}(Y)=\operatornamewithlimits{\textstyle\bigoplus}_{i=1}^{r}\mathcal{U}_{K}(Y)(\operatorname{codim}b_{i})$.
@@
${\overline{\mathop{\mathrm{CH}}\nolimits}^{*}_{K}(X\times Y)\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\mathop{\mathrm{CH}}\nolimits^{*}_{K}(\overline{X}\times\overline{Y})\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\mathop{\mathrm{CH}}\nolimits^{*}_{K}(\overline{X}\times\overline{Y})/I_{X\times Y}}$${\overline{\mathop{\mathrm{CH}}\nolimits}^{*}_{K}(X)\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\mathop{\mathrm{CH}}\nolimits^{*}_{K}(\overline{X})\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\mathop{\mathrm{CH}}\nolimits^{*}_{K}(\overline{X})/I_{X}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$
@@
${\mathop{\mathrm{CH}}\nolimits^{*}_{K}(\overline{X}\times\overline{Y},\overline{\pi}^{\prime})\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\mathop{\mathrm{CH}}\nolimits^{*}_{K}(\overline{X}\times\overline{Y})\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\mathop{\mathrm{CH}}\nolimits^{*}_{K}(\overline{X}\times\overline{Y})/I_{X\times Y}}$${\mathop{\mathrm{CH}}\nolimits^{*}_{K}(\overline{X},\overline{\pi})\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\mathop{\mathrm{CH}}\nolimits^{*}_{K}(\overline{X})\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\mathop{\mathrm{CH}}\nolimits^{*}_{K}(\overline{X})/I_{X}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$
@@ 8. $J$-invariant for groups of outer type
${H^{1}(F,H)\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{R_{H}}$${H^{3}(F,\mathbb{Q}/\mathbb{Z}(2))\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{n\cdot}$${H^{1}(F,G)\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{R_{G}}$${H^{3}(F,\mathbb{Q}/\mathbb{Z}(2)).}$
@@
[…]
@@
Let $\mathfrak{X}$ be a generically quasi-split variety under the action of the group $\operatorname{Aut}_{K}(B,\tau)$. Let $U$ be the normed upper motive of $\mathfrak{X}$. Consider the variety $\mathcal{Y}:=\mathfrak{X}\times R_{K/F}(\mathop{\mathrm{SB}}(A))$}. The variety $\mathcal{Y}$ is generically quasi-split under the action of the group
$\operatorname{Aut}_{K}(B,\tau)\times R_{K/F}(\operatorname{\mathrm{PGL}}_{1}(A))$}.
Since both varieties $\mathfrak{X}_{F(\mathcal{Y})}$ and $\mathcal{Y}_{F(\mathfrak{X})}$ have rational points, the normed upper motives of $\mathfrak{X}$ and $\mathcal{Y}$ are isomorphic: $\mathcal{U}_{K}(\mathcal{Y})\simeq\mathcal{U}_{K}(\mathfrak{X})=:U$. Denote by $m$ and $F_{A}$ respectively the degree of $A$ and the function field of the variety $R_{K/F}(\mathop{\mathrm{SB}}(A))$.
@@
Our next goal is to describe the quotient $P(\overline{\mathop{\mathrm{CH}}\nolimits}_{K}^{*}(\mathcal{Y}),t)/P(\overline{\mathop{\mathrm{CH}}\nolimits}_{K}^{*}(\mathfrak{X}_{F_{A}}),t)$. We follow similar arguments as in [Zh24, §3]. We refer to Section 3 for the notation $R(h)$ below.
@@
Since both varieties $\mathfrak{X}$ and $R_{K/F}(\mathop{\mathrm{SB}}(A))$ are quasi-split over the function field of $\mathfrak{X}$, we can follow the proof of [Zh24, Lemma 3.2] by replacing the usual Chow rings $\mathop{\mathrm{CH}}\nolimits$ by the normed Chow rings $\mathop{\mathrm{CH}}\nolimits_{K}$. ∎
Using Lemma 8.6, similarly as in [Zh24, Proposition 3.4] we can show that the polynomial $P(\overline{\mathop{\mathrm{CH}}\nolimits}_{K}^{*}(\mathfrak{X}_{F_{A}}),t)$ divides $P(\overline{\mathop{\mathrm{CH}}\nolimits}_{K}^{*}(\mathcal{Y}),t)$ and the quotient is given by $F(t):=\operatornamewithlimits{\textstyle\sum}_{i\in J}t^{i}$, where
(8.7) $J=\{0\leq j<m\mid 1\times R(h)^{j}\text{ is }F\text{-rational in }\mathop{\mathrm{CH}}\nolimits_{K}^{*}(\overline{\mathcal{Y}})\}\,.$
@@
We show next that the parameter $k_{1}$ of the $J$-invariant satisfies $k_{1}<v_{2}(\deg B)$. We assume that we are in type $^{2}\mathrm{A}_{2n+1}$ and, thus, $\deg B=2n+2$. Consider the generically quasi-split variety $Y$ of parabolic subgroups of the group $\operatorname{Aut}_{K}(B,\tau)$ of type
$\{1,2,\ldots,2n+1\}\setminus\{1,n+1,2n+1\}.$
This variety is generically quasi-split by [Ti66, p. 55, type $^{2}\mathrm{A}_{n}$]. Then the normed Poincaré polynomial of $Y$ must be divisible by the Poincaré polynomial of the normed upper motive $U$ of the variety of Borel subgroups. Thus, the polynomial
@@
[…]
Note that the variety $\mathfrak{X}$ possesses a rational point over the function field of the variety $\mathcal{Z}\times R_{K/F}(\mathop{\mathrm{SB}}(A))$. Therefore, by [Zh24, Lemma 2.3] the cycle
@@
is $F$-rational. It follows that $R(h)^{j}\in\mathop{\mathrm{CH}}\nolimits_{K}^{*}(\overline{R_{K/F}(\mathop{\mathrm{SB}}(A))})$} is rational over $F(\mathcal{Z})$. Therefore, passing to $K$ we find a rational cycle over $K(\mathcal{Z})$ of the form
$h^{j}\times h^{j}+\alpha\in\mathop{\mathrm{Ch}}\nolimits^{*}(\overline{\mathop{\mathrm{SB}}(A)\times\mathop{\mathrm{SB}}(A)})$}
in the ordinary Chow group, where $\alpha$ is some cycle which goes to zero in the respective normed Chow group.
@@
In the notation of Proposition 8.2 ($\deg B$ is even) we have: $j_{2}=0$ or 11, and $j_{2}=0$ iff the discriminant algebra $D(B,\tau)$ (see [Ta76], [KMRT, Definition (10.28)]) is split.
@@
In the notation of Proposition 8.2 ($\deg B=2n+2$ is even) let
$(j_{1},j_{2},j_{3},\ldots,j_{n+2})$
be the $J$-invariant of $(B,\tau)$. Then the $J$-invariant of $(B,\tau)_{F(\mathop{\mathrm{SB}}(D(B,\tau)))}$ equals
$(j_{1},0,j_{3},\ldots,j_{n+2}),$
i.e., apart from the value of $j_{2}$ the $J$-invariant does not change.
@@
The proof is the same as of [Zh24, Theorem 4.1(1)]. ∎
@@
By [Ka12a, Theorem 5.1] the ordinary Chow motive modulo $2$ of the generic orthogonal Grassmannian $\mathrm{OGr}$ of isotropic (n1)(n-1)-dimensional subspaces (it corresponds to the parabolic subgroup of type $\{1,2,\ldots,n-2\}$) is indecomposable and the same method of the proof shows that its normed motive is indecomposable as well. Since for a split algebra $A$ the upper motive of this Grassmannian is isomorphic to the upper motive of the variety of Borel subgroups, its Poincaré polynomial equals the Poincaré polynomial of $^{2}\mathrm{D}_{n}/P_{\{1,2,\ldots,n-2\}}$. The formula for the latter Poincaré polynomial follows directly from Proposition 6.1, namely, it equals
@@
Since the motive of the generic orthogonal Grassmannian of isotropic (n1)(n-1)-dimensional subspaces is indecomposable, it follows that the ring $C^{*}$ from sequence (5.3) equals in the $^{2}\mathrm{D}_{n}$ case the normed Chow ring $\mathop{\mathrm{CH}}\nolimits^{*}_{K}(\mathrm{OGr}(n-1,q))$ for a quasi-split quadratic form $q$ of dimension $2n$.
@@
The (normed) $J$-invariant of a central simple algebra with an orthogonal involution $(A,\sigma)$ of a non-trivial discriminant is the concatenation of the $J$-invariant of the quadratic form associated with $\sigma_{F(\mathop{\mathrm{SB}}(A))}$ (computed above) and
$j_{1}=\min\{k_{1},v_{2}(\mathop{\mathrm{ind}}A)\}$}.
@@
Let $U$ denote the normed upper motive of a generically quasi-split variety for the group $\operatorname{\mathrm{PGO}}^{+}(A,\sigma)$ and $U^{F_{A}}$ the normed upper motive of the same variety over the function field $F_{A}$ of the Severi–Brauer variety $\mathop{\mathrm{SB}}(A)$. Proceeding as in [Zh24, §3] we arrive at the following identity for the Poincaré polynomials
@@
We show now that $2^{k_{1}}\mid(\frac{\deg A}{2}-1)=n-1$. Consider the generically quasi-split variety $Y$ of parabolic subgroups of type $\{1,2,\ldots,n\}\setminus\{1,n-1,n\}$. This variety is generically quasi-split by [Ti66, p. 57, type $^{2}\mathrm{D}_{n}$]. Then the normed Poincaré polynomial of $Y$ must be divisible by the Poincaré polynomial of the normed upper motive $U$ of the variety of Borel subgroups. Thus, the polynomial
@@
Computing the quotient $\frac{P(\mathop{\mathrm{CH}}\nolimits^{*}_{K}(Y),t)}{P(\mathop{\mathrm{CH}}\nolimits^{*}_{K}(U),t)}$ we obtain $\frac{t^{n-1}-1}{t^{2^{k_{1}}}-1}$ which is a polynomial iff $2^{k_{1}}$ divides n1n-1.
@@
The normed $J$-invariant of the respective group of type $^{2}\mathrm{D}$ equals $(j_{1},j_{2},\ldots,j_{r})$, where $j_{i}$ of degree $d_{i}$ equals $1$, if $d_{i}=\dim\pi_{s-1}/2-1$ and equals 0 otherwise.
@@
where $\ldots$" stands for a sum of monomials of degree $\geq 3$.
Since $A$ is split, note that the generator $w$ of $\mathop{\mathrm{CH}}\nolimits^{1}_{K}(\overline{X})\simeq\mathbb{F}_{2}$ is $F$-rational, and moreover $w^{2}\neq 0$. Hence, we have $P(\overline{\mathop{\mathrm{CH}}\nolimits}_{K}^{*}(X),t)=1+t+t^{2}+\ldots\,$, where $\ldots$" stands for a sum of monomials of degree $\geq 3$.